OpenStudy (anonymous):

simplify ((x^2 + 3)-(-5^2 + 3)) / (x--5)

OpenStudy (saifoo.khan):

where are you stuck in that?

OpenStudy (anonymous):

(x^2+3-(-25+3))/(x-5)= (x^2+3+22)/(x-5)= (x^2+25)/(x-5)